OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x: x f(x) = 15x - 12 g(x) = 3x -2 0.11 -1 -27 0.33 0 -12 1 2 Complete the chart and determine the solution of the equation f(x) = g(x) x = -1 x = 0 x = 1 x = 3

OpenStudy (anonymous):

Um, can you fix the table? It's hard to red.

OpenStudy (anonymous):

*read

OpenStudy (anonymous):

x f(x) = 15x - 12 g(x) = 3x -2 0.11 -1 -27 0.33 0 -12 1 2

OpenStudy (anonymous):

it does look confusing im sorry

OpenStudy (anonymous):

Well, for the second part, we write: 15x - 12 = 3x Can you solve that?

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

\(15x - 12 = 3x\) Add 12 to both sides: \(15x = 3x + 12\) Subtract 3x to both sides: \(12x = 12\) Divide 12 to both sides: \(x = 1~\Large\color{lime} \checkmark\) You got it!

OpenStudy (anonymous):

awesome! thanks!

OpenStudy (anonymous):

No problem.